# https://xiaojiu-z.github.io/Stable-Makeup.github.io/

### v1

@@
Abstract
Current makeup transfer methods are limited to simple makeup styles, making them difficult to apply in real-world scenarios. In this paper, we introduce Stable-Makeup, a novel diffusion-based makeup transfer method capable of robustly transferring a wide range of real-world makeup, onto user-provided faces. Stable-Makeup is based on a pre-trained diffusion model and utilizes a Detail-Preserving (D-P) makeup encoder to encode makeup details. It also employs content and structural control modules to preserve the content and structural information of the source image. With the aid of our newly added makeup cross-attention layers in U-Net, we can accurately transfer the detailed makeup to the corresponding position in the source image. After content-structure decoupling training, Stable-Makeup can maintain content and the facial structure of the source image. Moreover, our method has demonstrated strong robustness and generalizability, making it applicable to various tasks such as cross-domain makeup transfer, makeup-guided text-to-image generation and so on. Extensive experiments have demonstrated that our approach delivers state-of-the-art (SOTA) results among existing makeup transfer methods and exhibits a highly promising with broad potential applications in various related fields.
Keywords:
Makeup Transfer, D-P Makeup Encoder, Diffusion
Figure 1: Our proposed framework, Stable-Makeup, is a novel diffusion-based method for makeup transfer that can robustly transfer a diverse range of real-world makeup styles, from light to extremely heavy makeup.
*** Work done during internship at Xiaohongshu Inc.
$\dagger$ Corresponding author
1 Introduction
@@
From the perspective of the existing technical route, current makeup transfer techniques [21, 17, 7, 26, 47, 12, 22, 43, 50, 40, 19, 51] rely heavily on Generative Adversarial Network (GAN) based approaches which have demonstrated potential in digital cosmetic applications. However, these approaches fall short when confronted with the diversity of real-world makeup styles, especially when translating high-detailed and creative cosmetics, such as those found in cosplay or movie character imitations, onto real faces. This limitation not only restricts their applicability but also hinders their effectiveness in accurately capturing the essence of personalized and intricate makeup designs. Recognizing this gap, our research introduces Stable-Makeup, a novel approach leveraging diffusion-based methodologies to transcend the boundaries of existing makeup transfer methods.
From a data perspective, existing makeup datasets lack diversity and cannot accommodate real-world makeup transfer. To address this data shortage, we propose an automatic data construction pipeline that employs large language model and generative model to edit real human face images and create paired before-and-after makeup images. Ultimately, our dataset comprises 20k image pairs, covering a wide range of makeup styles from light to heavy makeup. This dataset has the potential to facilitate research in the makeup transfer field and enable the development of more robust and accurate models.
Our Stable-Makeup, which is built on a pre-trained diffusion model, and comprises three key components: the Detail-Preserving Makeup Encoder, Makeup Cross-attention Layers and the Content and Structural Control Modules. To preserve makeup details, we employ a multi-layer strategy in our D-P makeup encoder to encode the makeup reference image into multi-scale and spatial-aware detail makeup embeddings. The content control module is designed to maintain pixel-level content consistency with source image. The structural control module is utilized to introduce facial structure, improving the consistency between the generated image and the facial structure of the source image. To achieve semantic alignment between the intermediate features of the Unet-encoded source image and the detail makeup embeddings, we extended the U-Net architecture by incorporating a makeup branch composed of cross-attention layers. Through our proposed content and structural decoupling training strategy, we can further maintain the facial structure of the source image. As exhibited in Fig. 1, By integrating a meticulously curated dataset alongside a comprehensive pipeline from data acquisition to model training, we lay the foundation for a robust framework capable of handling the nuanced dynamics of makeup transfer with unprecedented precision and adaptability.
@@
  • In this paper, we present Stable-Makeup, the first diffusion-based makeup transfer method to the best of our knowledge. Our experimental results demonstrate state-of-the-art performance over existing makeup transfer methods. Moreover, Stable-Makeup demonstrates high robustness and generalizability, making it applicable to various tasks.
@@
  • We designed an automatic pipeline to create a high-quality, diverse paired dataset of makeup before and after images. After content-structure decoupling training on this dataset, we can further maintain the content and structure of the source image.
2 Related Works
2.1 Facial Makeup Transfer
Traditional makeup transfer methods employ image processing techniques such as facial landmark extraction and detection[49, 41]. Meanwhile, contemporary deep learning-based makeup transfer methods demonstrate high robustness and generalizability to diverse makeup styles, Generative Adversarial Networks (GANs)[58, 5, 21, 17, 7, 26, 47, 12, 22, 43, 50, 40, 19, 15, 51] as an example, have been extensively applied in facial makeup synthesis tasks.
In the domain of image transfer, Beauty-GAN[21] driven by pixel-level Histogram Matching and utilizes an array of loss functions to train the primary network. PSGAN[17] addresses the challenge of transferring makeup between images with significant facial expression variations and focuses on specific facial regions for transfer. To expand the scope of makeup transfer tasks beyond mere color transfer in specific facial areas, CPM[26] introduces pattern addition into the makeup transfer process. Furthermore, SCGAN[7] leverages a Part-specific Style Encoder to decouple makeup styles on a component-wise basis. Additionally, RamGAN[47] focuses on component consistency by integrating a region-attentive morphing module.
@@ 2.1 Facial Makeup Transfer
2.2 Diffusion Models
Recently, diffusion models have ushered in significant advancements in multi-modal image generation tasks, encompassing text-to-image generation[30, 27, 34, 31, 35], image editing[6, 20, 9, 3, 48, 24, 2, 56, 42], controllable generation[53, 25, 57, 23] and subject-driven generation[55, 14, 1, 16, 10, 33, 32]. Methods based on diffusion paradigm, along with their subsequent adaptations, have demonstrated state-of-the-art results in various image generation tasks.
The diffusion model was initially introduced based on the physical inverse diffuse process [37]. Subsequently, DDPM[13] demonstrated that this approach is viable for image generation by denoising a Gaussian-noised image. DDIM[38] introduced a faster non-Markovian sampling process and an implicit modeling of the denoising step, both of which significantly enhance the model’s efficiency. Latent Diffusion Models[31] utilize an Autoencoder to encode images into latent embeddings, effectively increasing computational efficiency while preserving the essential features of the input image. In recent years, there have been significant advancements in diffusion-based text-to-image generation. For instance, Imagen[34] and Stable Diffusion[31] have trained diffusion models on larger datasets, making them the mainstream for generating large images. DeepFloyd IF[35] has utilized a triple-cascade diffusion model, which has significantly enhanced the text-to-image generation capability, even generating correct fonts. Stable Diffusion XL[27], a two-stage cascade diffusion model, which has greatly improved the generation of high-frequency details, small object features, and overall image color. As for controllable generation, ControlNet[54] employs a bypass structure featuring a trainable copy network as the input mechanism for control conditions, effectively integrating a variety of control conditions such as Canny Edge, Depth, and Human Pose, etc. As for subject generation, DreamBooth[32] Dreambooth trains a separate Unet for each subject, resulting in excellent consistency of character. However, the high computational cost limits its practical application. Meanwhile, diffusion based image editing methods have also demonstrated remarkable progress. For example, Ledits[42] introduces a training-free image manipulation method, utilizing an inversion approach for processing.
@@ 2.2 Diffusion Models
3 Methodology
In this part, we introduce the detailed design of our Stable-Makeup in Section 3.1. Then, the specifics of our training and inference process are detailed in Section 3.2. Preliminaries about diffusion models [31], Controlnet [54] and CLIP [29] can be found in supplementary.
3.1 Stable-Makeup
3.1.1 Overview.
The overall methodology is illustrated in Fig. 2. Formally, the target of Stable-Makeup is to transfer the reference makeup from a given image $\mathit{I_{m}}$ onto a user-provided source image $\mathit{I_{s}}$, resulting in a target image $\mathit{I_{t}}$ that exhibits the desired makeup style. Therefore, Stable-Makeup necessitates the simultaneous transfer of fine-grained makeup information and preservation of structural information. To achieve this, Stable-Makeup first utilizes a Detail-Preserving Makeup Encoder to extract the multi-scale and spatial-aware features of the reference makeup as detailed makeup embeddings $\mathit{E_{m}}$. Next, content encoder and structural encoder are employed to encode the source image as content features $\mathit{F_{c}}$ and the facial structure control image $\mathit{I_{c}}$ (obtained from an off-the-shelf face parsing method [28].) as structure features $\mathit{F_{s}}$, respectively. These two features are then incorporated together and fed into the U-Net. The detailed makeup embeddings $\mathit{E_{m}}$ along with the content features $\mathit{F_{c}}$ and structure features $\mathit{F_{s}}$ are integrated into the diffusion U-Net as conditions to guide the image generation process. During target image $\mathit{I_{t}}$ generation, Stable-Makeup uses newly added cross-attention layers to implicitly align the multi-scale detailed makeup embeddings $\mathit{E_{m}}$ with the intermediate feature maps of facial region in source image at different layers in U-Net. The alignment ensures that the generated target image contains both the detailed information of the reference makeup and maintains the content and structure of the source image $\mathit{I_{s}}$. With our Content and Structure Decoupling Training, Stable-Makeup can further maintain the content and facial structure of the source image.
Figure 2: Overall schematics of our method. Given a source image $\mathit{I_{s}}$, a reference makeup image $\mathit{I_{m}}$ and an obtained facial structure control image $\mathit{I_{c}}$, Stable-Makeup utilizes D-P makeup encoder to encode $\mathit{I_{m}}$. Content and structural encoders are used to encode $\mathit{I_{s}}$ and $\mathit{I_{c}}$ respectively. With the aid of the makeup cross-attention layers, Stable-Makeup aligns the facial regions of $\mathit{I_{s}}$ and $\mathit{I_{m}}$, enabling successful transfers the intricate makeup details. After content-structure decoupling training, Stable-Makeup further maintains content and structure of $\mathit{I_{s}}$.
3.1.2 Detail-Preserving Makeup Encoder.
@@ 3.1 Stable-Makeup
Notably, some previous works[52, 46, 11] utilize linear layer mapping of CLIP features. This design cannot preserve the spatial information between input features. Different from them, we utilize a self-attention layer to map the multi-layer CLIP embeddings (256-dimensional patch features and 1-dimensional CLS features) of the reference makeup image and concatenate on feature dimension. This approach effectively preserves the spatial, multi-scale, and global information of the makeup image, thereby facilitating improved alignment of intermediate related features within the U-Net architecture. Consequently, it yields more precise preservation of fine-grained details and enhanced transfer outcomes. Combining both spatial and scale feature maps, our D-P makeup encoder yields the most complete information for makeup. Furthermore, we provide additional comparative analysis in supplementary to demonstrate the superiority of our method.
3.1.3 Makeup Cross-attention Layers.
The goal of the makeup cross-attention layers is to integrate multi-scale detailed makeup embeddings into the U-Net and align them appropriately with the intermediate feature maps of facial region in the source image, enabling the model to accurately transfer the makeup onto the facial region in the source image. Specifically, as shown in the figure 2, in each transformer block of the U-Net, we retain the original text cross-attention and add Makeup Cross-attention Layers. The multi-scale detailed makeup embeddings $\mathit{E_{m}}$ are input to the makeup cross-attention layers, serves as the K(key) and V(value) features. The makeup cross-attention layers and the text cross-attention layers in the original U-Net share the Q feature. With the help of the cross-attention mechanism, the intermediate facial features of source image and makeup image are well-aligned in the stable diffusion U-Net. Finally, we simply add the output of makeup cross-attention to the output of text cross-attention.
@@
3.1.4 Content and Structural Control Modules.
To maintain the structural consistency of the source image during makeup transfer, we utilize Content and Structural Control Modules, which consists of two adapted ControlNets serving as encoders. The target of the content encoder is to maintain content consistency and consistency in non-facial areas of the original image, and it uses the ControlNet to encode the source image. To control facial structures during the makeup transfer process, we added an extra structural encoder. In order to increase spatial constraints and introduce more facial information, such as facial shape or mouth closure, we utilized dense lines of varying colors that were drawn based on the different locations of facial key points. These lines served as conditions for the makeup transfer process, enabling us to achieve great control over the resulting facial structures. We use the text embedding of empty text as the cross-attention input for the two ControlNet encoders. Finally, we directly add the outputs of the content encoder and structural encoder to the Stable Diffusion model.
3.2 Model Training and Inference
3.2.1 Makeup-real Dataset.
As depicted in Fig. 3, Ours pipeline to create a large-scale makeup pairing dataset (Makeup-real dataset) by leveraging the capabilities of LLM and pre-trained diffusion involves three main steps: Firstly, we utilize GPT4 to generate 1000 prompts using the template "Make it {} makeup" to cover a wide range of makeup styles. After filtering out duplicate words, we obtain 300 different prompts. Secondly, we sample 20,000 faces from the FFHQ[18] dataset as non-makeup images and apply makeup to these images using existing text-based real image editing algorithm [42]. For each face image, we randomly select a makeup style from the 300 prompts obtained in the first step, resulting in 20,000 face images with consistent identities and makeup styles. Finally, to prevent the unwanted effects of editing non-face regions during the edit process, we use a face segmentation model to separate the face region from the background and other non-face regions. We composite the non-face regions of the original face images onto the edited face images to obtain the final makeup pairing dataset. [Uncaptioned image] Figure 3: The procedure for training data generation. Finally, our dataset includes 20,000 image pairs covering a wide range of makeup styles from light to heavy makeup, represents a significant improvement over previous datasets in terms of diversity and quantity. In our supplementary materials, we also present additional visual results for the alternative methods in Step 2.
3.2.2 Content-Structure Decoupling Training.

### v2

# Stable-Makeup: When Real-World Makeup Transfer Meets Diffusion Model

@@
Abstract.
Current makeup transfer methods are limited to simple makeup styles, making them difficult to apply in real-world scenarios. In this paper, we introduce Stable-Makeup, a novel diffusion-based makeup transfer method capable of robustly transferring a wide range of real-world makeup, onto user-provided faces. Stable-Makeup is based on a pre-trained diffusion model and utilizes a Detail-Preserving (D-P) makeup encoder to encode makeup details. It also employs content and structural control modules to preserve the content and structural information of the source image. With the aid of our newly added makeup cross-attention layers in U-Net, we can accurately transfer the detailed makeup to the corresponding position in the source image. After content-structure decoupling training, Stable-Makeup can maintain the content and the facial structure of the source image. Moreover, our method has demonstrated strong robustness and generalizability, making it applicable to various tasks such as cross-domain makeup transfer, makeup-guided text-to-image generation, and so on. Extensive experiments have demonstrated that our approach delivers state-of-the-art results among existing makeup transfer methods and exhibits a highly promising with broad potential applications in various related fields. Code released: https://github.com/Xiaojiu-z/Stable-Makeup
Diffusion Model, Makeup Transfer, Image Generation
isbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Computer vision
Figure 1. Our proposed framework, Stable-Makeup, is a novel diffusion-based method for makeup transfer that can robustly transfer a diverse range of real-world makeup styles, from light to extremely heavy makeup.
[…]
1. Introduction
@@
From the perspective of the existing technical route, current makeup transfer techniques (Li et al., 2018; Jiang et al., 2020; Deng et al., 2021; Nguyen et al., 2021; Xiang et al., 2022; Gu et al., 2019; Liu et al., 2021; Wan et al., 2022; Yan et al., 2023; Sun et al., 2023; Kips et al., 2020; Yang et al., 2022) rely heavily on Generative Adversarial Network (GAN) based approaches which have demonstrated potential in digital cosmetic applications. However, these approaches fall short when confronted with the diversity of real-world makeup styles, especially when translating high-detailed and creative cosmetics, such as those found in cosplay or movie character imitations, onto real faces. This limitation not only restricts their applicability but also hinders their effectiveness in accurately capturing the essence of personalized and intricate makeup designs. Recognizing this gap, our research introduces Stable-Makeup, a novel approach leveraging diffusion-based methodologies to transcend the boundaries of existing makeup transfer methods.
From a data perspective, existing makeup datasets lack diversity and cannot accommodate real-world makeup transfer. To address this data shortage, we propose an automatic data construction pipeline that employs a large language model and generative model to edit real human face images and create paired before-and-after makeup images. Ultimately, our training data comprises 20k image pairs, covering a wide range of makeup styles from light to heavy makeup. This automatic data construction pipeline has the potential to facilitate research in the makeup transfer field and enable the development of more robust and accurate models.
Our Stable-Makeup, which is built on a pre-trained diffusion model, comprises three key components: the Detail-Preserving Makeup Encoder, Makeup Cross-attention Layers, and the Content and Structural Control Modules. To preserve makeup details, we employ a multi-layer strategy in our D-P makeup encoder to encode the makeup reference image into multi-scale detail makeup embeddings. The content control module is designed to maintain pixel-level content consistency with the source image. The structural control module is utilized to introduce facial structure, improving the consistency between the generated image and the facial structure of the source image. To achieve semantic alignment between the intermediate features of the Unet-encoded source image and the detail makeup embeddings, we extended the U-Net architecture by incorporating a makeup branch composed of cross-attention layers. Through our proposed content and structural decoupling training strategy, we can further maintain the facial structure of the source image. As exhibited in Fig. 1, By integrating a meticulously curated dataset alongside a comprehensive pipeline from data acquisition to model training, we lay the foundation for a robust framework capable of handling the nuanced dynamics of makeup transfer with unprecedented precision and adaptability.
@@
  • Stable-Makeup is the first diffusion-based makeup transfer method to the best of our knowledge. Our experimental results demonstrate state-of-the-art performance over existing makeup transfer methods. Moreover, Stable-Makeup demonstrates high robustness and generalizability, making it applicable to various tasks.
@@
  • We designed an automatic pipeline to create a high-quality, diverse paired dataset of makeup before and after images. After content-structure decoupling training on our training data, we can further maintain the content and structure of the source image.
2. Related Works
2.1. Facial Makeup Transfer
Traditional makeup transfer methods employ image processing techniques such as facial landmark extraction and detection(Xu et al., 2013; Tong et al., 2007). Meanwhile, contemporary deep learning-based makeup transfer methods demonstrate high robustness and generalizability to diverse makeup styles, Generative Adversarial Networks (GANs)(Zhu et al., 2017; Chang et al., 2018; Li et al., 2018; Jiang et al., 2020; Deng et al., 2021; Nguyen et al., 2021; Xiang et al., 2022; Gu et al., 2019; Liu et al., 2021; Wan et al., 2022; Yan et al., 2023; Sun et al., 2023; Kips et al., 2020; Hu et al., 2022a; Yang et al., 2022) as an example, have been extensively applied in facial makeup synthesis tasks.
In the domain of image transfer, Beauty-GAN(Li et al., 2018) is driven by pixel-level Histogram Matching and utilizes an array of loss functions to train the primary network. However, its effectiveness is limited to frontal facial images, posing a challenge when transferring makeup between images with notable facial pose variations. PSGAN(Jiang et al., 2020) addresses the challenge of transferring makeup between images with significant facial expression variations and focuses on specific facial regions for transfer. To expand the scope of makeup transfer tasks beyond mere color transfer in specific facial areas, CPM(Nguyen et al., 2021) introduces pattern addition into the makeup transfer process. Furthermore, SCGAN(Deng et al., 2021) leverages a Part-specific Style Encoder to decouple makeup styles on a component-wise basis. Additionally, RamGAN(Xiang et al., 2022) focuses on component consistency by integrating a region-attentive morphing module. Although these GAN-based methods have achieved significant success, they exhibit limitations in transferring complex makeup styles in real-world scenarios. Moreover, the majority of these approaches necessitate the alignment of facial regions, further constraining their practical applicability.
@@ 2.1. Facial Makeup Transfer
2.2. Diffusion Models
Recently, diffusion models have ushered in significant advancements in multi-modal image generation tasks, encompassing text-to-image generation(Ramesh et al., 2022; Podell et al., 2023; Saharia et al., 2022a; Rombach et al., 2022; Saharia et al., 2022b), image editing(Chefer et al., 2023; Li et al., 2023; Epstein et al., 2023; Cao et al., 2023; Xie et al., 2023; Mou et al., 2023a; Bodur et al., 2023; Zhang et al., 2023a; Tsaban and Passos, 2023), controllable generation(Zhang and Agrawala, 2023; Mou et al., 2023b; Zhao et al., 2023; Ma et al., 2023) and subject-driven generation(Zhang et al., 2024; Hu et al., 2022b; Arar et al., 2023; Jia et al., 2023; Gal et al., 2022; Ruiz et al., 2023, 2022). Methods based on the diffusion paradigm, along with their subsequent adaptations, have demonstrated state-of-the-art results in various image generation tasks.
As for controllable generation, ControlNet(Zhang et al., 2023b) employs a bypass structure featuring a trainable copy network as the input mechanism for control conditions, effectively integrating a variety of control conditions such as Canny Edge, Depth, and Human Pose, etc. As for subject generation, DreamBooth(Ruiz et al., 2022) Dreambooth trains a separate Unet for each subject, resulting in excellent consistency of character. Meanwhile, diffusion-based image editing methods have also demonstrated remarkable progress. For example, Ledits(Tsaban and Passos, 2023) introduces a training-free image manipulation method, utilizing an inversion approach for processing.
@@ 2.2. Diffusion Models
3. Methodology
In this part, we introduce the detailed design of our Stable-Makeup. Then, the specifics of our training and inference process are in a detailed description. Preliminaries about diffusion models (Rombach et al., 2022), Controlnet (Zhang et al., 2023b) and CLIP (Radford et al., 2021) can be found in supplementary.
3.1. Stable-Makeup
3.1.1. Overview.
The overall methodology is illustrated in Fig. 2. Formally, the target of Stable-Makeup is to transfer the reference makeup from a given image $\mathit{I_{m}}$ onto a user-provided source image $\mathit{I_{s}}$, resulting in a target image $\mathit{I_{t}}$ that exhibits the desired makeup style. Therefore, Stable-Makeup necessitates the simultaneous transfer of fine-grained makeup information and the preservation of structural information. To achieve this, Stable-Makeup first utilizes a Detail-Preserving Makeup Encoder to extract the multi-scale features of the reference makeup as detailed makeup embeddings $\mathit{E_{m}}$. Next, content encoder and structural encoder are employed to encode the source image as content features $\mathit{F_{c}}$ and the facial structure control image $\mathit{I_{c}}$ (obtained from an off-the-shelf face parsing method (Prados-Torreblanca et al., 2022).) as structure features $\mathit{F_{s}}$, respectively. These two features are then incorporated together and fed into the U-Net. The multi-scale detailed makeup embeddings $\mathit{E_{m}}$ along with the content features $\mathit{F_{c}}$ and structure features $\mathit{F_{s}}$ are integrated into the diffusion U-Net as conditions to guide the image generation process. During target image $\mathit{I_{t}}$ generation, Stable-Makeup uses newly added cross-attention layers to implicitly align the multi-scale detailed makeup embeddings $\mathit{E_{m}}$ with the intermediate feature maps of facial region in source image at different layers in U-Net. The alignment ensures that the generated target image contains both the detailed information of the reference makeup and maintains the content and structure of the source image $\mathit{I_{s}}$. With our Content and Structure Decoupling Training, Stable-Makeup can further maintain the content and facial structure of the source image.
Figure 2. Overall schematics of our method. Given a source image $\mathit{I_{s}}$, a reference makeup image $\mathit{I_{m}}$ and an obtained facial structure control image $\mathit{I_{c}}$, Stable-Makeup utilizes the D-P makeup encoder to encode $\mathit{I_{m}}$. Content and structural encoders are used to encode $\mathit{I_{s}}$ and $\mathit{I_{c}}$ respectively. With the aid of the makeup cross-attention layers, Stable-Makeup aligns the facial regions of $\mathit{I_{s}}$ and $\mathit{I_{m}}$, enabling successful transfers of the intricate makeup details. After content-structure decoupling training, Stable-Makeup further maintains content and structure of $\mathit{I_{s}}$.
3.1.2. Detail-Preserving Makeup Encoder.
@@ 3.1. Stable-Makeup
Notably, some previous works(Ye et al., 2023; Wei et al., 2023; Gal et al., 2023) utilize linear layer mapping of CLIP features. This design cannot preserve the spatial information between input features. Different from them, we utilize a self-attention layer to map the multi-layer CLIP embeddings (256-dimensional patch features and 1-dimensional CLS features) of the reference makeup image and concatenate on feature dimension. This approach effectively preserves the spatial, multi-scale, and global information of the makeup image, thereby facilitating improved alignment of intermediate-related features within the U-Net architecture. Consequently, it yields more precise preservation of fine-grained details and enhanced transfer outcomes. Combining both spatial and scale feature maps, our D-P makeup encoder yields the most complete information for makeup. Furthermore, we provide additional comparative analysis in supplementary to demonstrate the superiority of our method.
3.1.3. Makeup Cross-attention Layers.
The goal of the makeup cross-attention layers is to integrate multi-scale detailed makeup embeddings into the U-Net and align them appropriately with the intermediate feature maps ofthe facial region in the source image, enabling the model to accurately transfer the makeup onto the facial region in the source image. Specifically, as shown in the figure 2, in each transformer block of the U-Net, we retain the original text cross-attention and add Makeup Cross-attention Layers. The multi-scale detailed makeup embeddings $\mathit{E_{m}}$ are input to the makeup cross-attention layers and serve as the K(key) and V(value) features. The makeup cross-attention layers and the text cross-attention layers in the original U-Net share the Q feature. With the help of the cross-attention mechanism, the intermediate facial features of the source image and makeup image are well-aligned in the stable diffusion U-Net. Finally, we simply add the output of makeup cross-attention to the output of text cross-attention.
@@
3.1.4. Content and Structural Control Modules.
To maintain the structural consistency of the source image during makeup transfer, we utilize Content and Structural Control Modules, which consist of two adapted ControlNets serving as encoders. The target of the content encoder is to maintain content consistency and consistency in non-facial areas of the original image, and it uses ControlNet to encode the source image. To control facial structures during the makeup transfer process, we added an extra structural encoder. In order to increase spatial constraints and introduce more facial information, such as facial shape or mouth closure, we utilized dense lines of varying colors that were drawn based on the different locations of facial key points. These lines served as conditions for the makeup transfer process, enabling us to achieve great control over the resulting facial structures. We use the text embedding of empty text as the cross-attention input for the two ControlNet encoders. Finally, we directly add the outputs of the content encoder and structural encoder to the Stable Diffusion model.
3.2. Model Training and Inference
3.2.1. Training data collection.
As depicted in Fig. 3, our pipeline to create a large-scale makeup pairing dataset by leveraging the capabilities of LLM and pre-trained diffusion involves three main steps: Firstly, we utilize GPT4 to generate 1000 prompts using the template “Make it {} makeup” to cover a wide range of makeup styles. After filtering out duplicate words, we obtain 300 different prompts. Secondly, we sample 20,000 faces from the FFHQ(Karras et al., 2019) dataset as non-makeup images and apply makeup to these images using existing text-based real image editing algorithm (Tsaban and Passos, 2023). For each face image, we randomly select a makeup style from the 300 prompts obtained in the first step, resulting in 20,000 face images with consistent identities and makeup styles. Finally, to prevent the unwanted effects of editing non-face regions during the editing process, we use a face segmentation model to separate the face region from the background and other non-face regions. We composite the non-face regions of the original face images onto the edited face images to obtain the final makeup pairing dataset.
Figure 3. The procedure for training data generation.
Finally, our training data includes 20,000 image pairs covering a wide range of makeup styles from light to heavy makeup, which represents a significant improvement over previous datasets in terms of diversity and quantity. In our supplementary materials, we also present additional visual results for the alternative methods in Step 2.
Figure 4. Qualitative comparison of different methods. Our results outperform other methods in terms of makeup detail transfer, ranging from light makeup to personalized heavy makeup.
3.2.2. Content-Structure Decoupling Training.